Wage Incentive Plan, A producer of a perishable product offers a wage incentive to drivers of its trucks. A standard delivery takes an average of 20 hours. Drivers are paid at the rate of $10 per hour up to a maximum of 20 hours. There is an incentive for drivers to make the trip in less (but not too much less) than 20 hours. For each hour under 20, the hourly wage increases by $2.50. (the $2.50 – per – hour increase in wages applies for fractions of hours. That is if a trip takes 19.5 hours, the hourly wage increases by 0.5 x $2.50, or $1.25.) Determine the function w = f(n), where w= the hourly wage (in dollars) and n equals the number of hours to complete the trip.


1
Expert's answer
2021-05-07T09:12:02-0400

Answer :-



"w=f(n) = \begin{cases}
 0 & x<0 \\
 (10+2.5)\lfloor{x}\rfloor &0\leq x <20\\
 200 &x\geq20
\end{cases}"